# Hardware-Accelerated Ray Tracing for Discrete and Continuous Collision Detection on GPUs

### v1

@@ Abstract
This paper presents a set of simple and intuitive robot collision detection algorithms that show substantial scaling improvements for high geometric complexity and large numbers of collision queries by leveraging hardware-accelerated ray tracing on GPUs. It is the first leveraging hardware-accelerated ray-tracing for direct volume mesh-to-mesh discrete collision detection and applying it to continuous collision detection. We introduce two methods: Ray-Traced Discrete-Pose Collision Detection for exact robot mesh to obstacle mesh collision detection, and Ray-Traced Continuous Collision Detection for robot sphere representation to obstacle mesh swept collision detection, using piecewise-linear or quadratic B-splines. For robot link meshes totaling 24k triangles and obstacle meshes of over 190k triangles, our methods were up to 3 times faster in batched discrete-pose queries than a state-of-the-art GPU-based method using a sphere robot representation. For the same obstacle mesh scene, our sphere-robot continuous collision detection was up to 9 times faster depending on trajectory batch size. We also performed a detailed measurement of the volume coverage accuracy of various sphere/mesh pose/path representations to provide insight into the tradeoffs between speed and accuracy of different robot collision detection methods.
@@ I Introduction
The second scaling issue is the large number of collision detection queries needed for planning algorithms, often on the order of thousands of pose checks in addition to sub-path/edge checks. Additionally, a robot performing complex tasks may perform hundreds of motion planning calls when evaluating candidate tasks and ways to perform each task. Parallelism from CPU multithreading or SIMD instructions (as in [5]) saturates very quickly. GPUs scale much better, able to execute simultaneous instructions from thousands of times more threads. However, GPU-based collision detection [6] is difficult to implement—high-performance BVH construction and traversal on the GPU is very complex—leading to few well-maintained software implementations.
@@
    Ray-Traced Discrete-Pose Collision Detection (RT-DCD): Performs simple and intuitive mesh-to-mesh pose collision detection up to 3x faster than GPU-based mesh-to-sphere counterpart cuRobo [9] and up to 14x faster than CPU-based FCL [10], e.g., checking 4,096 articulated-arm poses against a $>$190k triangle mesh collision scene in 8 ms.
@@
    Ray-Traced Continuous Collision Detection (RT-CCD): Constructs swept volumes of sphere-approximated robots along trajectories using OptiX curves and detects collisions with mesh obstacles. Performs up to 9x faster than cuRobo, depending on trajectory batch size and curve type (piecewise linear vs. quadratic B-spline).
@@ II Background and Related Works
Collision-checking becomes more time-consuming as the geometric complexity of the robot and obstacles increases. It is quite common to simplify geometric representations of the robot and/or obstacles (e.g., using primitive shapes such as spheres or capsules), but this may be too conservative in cluttered scenes or tight spaces. Broad-phase/narrow-phase methods can also improve computational scaling through quick elimination of groups of collision geometries via different types of bounding volume hierarchies (BVH), such as AABB [1], OBB [2], SSV [3], and k-DOP [4].
@@ II-A CPU-Based DCD Methods
The Flexible Collision Library (FCL) [10] collision checks in 3D space with BVH acceleration and has seen perhaps the most widespread adoption in robotics, valued for its support for many geometry, collision pair, and query types. VAMP [5] further accelerated collision detection of sphere robots through CPU SIMD instructions. However, these libraries are CPU-bound, and CPU multithreading, SIMD, and BVH scaling are often not enough to handle both the geometric complexity and quantity of collision queries needed for low-conservatism real-time planning.
@@ II-B GPU-Based DCD Methods without Ray Tracing
Some recent methods like PairwiseNet [14] and ClearanceNet [15] leverage neural networks and GPUs for higher throughput, but the ClearanceNet could only achieve 91-96% detection accuracy and PairwiseNet requires training on each mesh pair for collision detection, whereas our DCD method detects all penetrating collisions between arbitrary meshes with minimal setup time.
@@ II-C CCD Methods
Conservative advancement (CA) methods [20, 21, 22] offer an alternative by computing a motion bound to determine the maximum forward sweep time step on a trajectory. The cuRobo library [9] uses a modified bisection and CA method on robots approximated by spheres, achieving CCD with excellent batch performance through parallel execution on GPUs. Our research proposes a new CCD method for sphere-approximated robots, using a GPU-based ray-tracing method to check against explicitly constructed swept volumes and achieving faster speeds depending on trajectory batch size and higher accuracy through swept curves which are higher-order than cuRobo’s method allows.
@@ II-D Hardware-Accelerated Ray Tracing
We have previously explored hardware-accelerated ray tracing for computing signed distances (which can serve as collision detection), finding contact points, and formulating trajectory optimization constraints between robot and obstacle meshes [7]. The Mochi engine [8] explored applications to DCD between spheres, triangle meshes, and generic geometries which can be AABB-bounded and have a point-containment function. However, its mesh DCD does not treat meshes as volumes (it cannot detect an obstacle fully contained in a robot mesh or vice-versa), and it has suboptimalities for common robot DCD applications (for example, not distinguishing desired robot/obstacle collision pairs and not exploring the best batching strategies for multi-query multi-link robot applications).
@@ III-A Ray Tracing Volume Mesh-to-Mesh Collision Detection
A further corner case is when the robot is fully contained inside an obstacle or vice versa, which is common in aerial, underwater, or space motion planning scenes with unanchored robots / obstacles. For these cases, it is sufficient to have watertight meshes with correctly oriented triangles (normals facing out) and to trace one maximum-length ray from the interior of each mesh toward all target meshes. If the ray hits more triangle back faces than front (which can be detected through OptiX), the ray origin is in collision. Note that this also handles target meshes which are not disjoint.
These tracing strategies produce exact volume mesh-to-mesh detection of non-penetrating collisions, which is sufficient for most robot motion planning applications.
@@ III-B Discrete-Pose Collision Detection
We developed three approaches to detect collisions between robot and obstacle volume meshes: ObsToRobot (Fig.LABEL:fig:_headline_d) traces rays along obstacles toward robot links, RobotToObs (Fig.LABEL:fig:_headline_r) traces rays along robot links posed by forward kinematics toward obstacles, and “two-way” simply combines the results of ObsToRobot and RobotToObs for exact detection results. Alg. 1 shows the inputs of each including a robot kinematics / link mesh model and a mesh obstacle scene. The output is a boolean array indicating whether each robot configuration is in collision with an obstacle. Each algorithm is GPU-parallelized to batch operations such as forward kinematics across robot configurations and aggregating results across rays, and each uses the method in Sec. III-A to identify in-collision mesh pairs.
@@
Data: Robot configurations $\{\theta_{i}\}$, robot model, scene, ${}_{\text{o}}$
1:$\bm{T_{f}}\leftarrow$ BatchForwardKinematics($\{\theta_{i}\}$, robot)
@@
To increase speedup, we precompute oriented bounding boxes (OBBs) for both obstacle and robot meshes to use for broad-phase collision detection (i.e. single-level BVHs for simplicity). These broad-phase results allow eliminating a large number of unnecessary posed links from IASs and rays from tracing steps. For example in the scene shown in Fig 2, only 1/4th of the associated OBBs are in collision. To aggregate useful posed links and rays from broadphase result masks, we use stream compaction [24] a parallel algorithm that gathers subsets of data of interest. Although stream compaction has poor scaling compared other common classes of parallel algorithms on the GPU, we found it was more than worth it to reduce inputs to ray-tracing steps, which tend to take by far the majority of compute time.
@@ III-C Continuous Collision Detection
For piecewise-linear swept volumes, we use batch forward kinematics to compute robot link sphere centers on poses along a path, linearly interpolating sphere centers between subsequent poses. For more accurate path approximations, we also implemented quadratic and cubic B-splines, for which we calculate the control points as follows.
For a B-spline with $n$ control points $\{\bm{p}_{i}\}$ and passing through $m$ points $\{\bm{q}_{j}\}$, we have the established relationship
$\underbrace{\begin{bmatrix}\bm{q}_{1}\\ \bm{q}_{2}\\ \vdots\\ \bm{q}_{m}\end{bmatrix}}_{\bm{Q}}=\underbrace{\begin{bmatrix}B_{0}(u_{1})&B_{1}(u_{1})&\dots&B_{n-1}(u_{1})\\ B_{0}(u_{2})&B_{1}(u_{2})&\dots&B_{n-1}(u_{2})\\ \vdots&\vdots&\ddots&\vdots\\ B_{0}(u_{m})&B_{1}(u_{m})&\dots&B_{n-1}(u_{m})\end{bmatrix}}_{\bm{\Phi}}\underbrace{\begin{bmatrix}\bm{p}_{1}\\ \bm{p}_{2}\\ \vdots\\ \bm{p}_{n}\end{bmatrix}}_{\bm{P}},\$ (1)
where $\{B_{i}\}$ are the basis functions for the B-spline [25] and $u_{j}$ is the curve parameter at $\bm{q}_{j}$. To minimize gaps between desired trajectory points $\bm{Q^{*}}=\{\bm{q^{*}}_{j}\}$ and actual curve points $\bm{Q}$, we use least squares:
$\bm{P}=\left(\bm{\Phi}^{\text{T}}\bm{\Phi}\right)^{-1}\bm{\Phi}^{\text{T}}\bm{Q^{*}}=\bm{\Phi}^{\dagger}\bm{Q^{*}}.$ (2)
The parameters $\{u_{i}\},\,i\in\{1,\dots,n\}$ are predetermined by the number of control points. Thus, the pseudo-inverse $\bm{\Phi}^{\dagger}$ remains constant, making each control point a simple constant-weighted sum of all trajectory points. Since $\bm{\Phi}^{\dagger}$ can then be reused for generating all curves across various links and poses, we store it in GPU constant memory for efficient broadcasting across instances of a parallel reduction kernel used to compute the curve control points.
One challenge is that OptiX meshes, spheres, and curves are all considered hollow, but curves are unique in that ray traces against them cannot detect back-face hits (intersections starting from a curve’s interior). To address this, we choose ray directions along the edges of each obstacle mesh such that they form a strongly-connected directed graph. This guarantees catching vertex collisions as any out-of-collision vertex must have a ray-directed path to any in-collision vertex. We precompute these graphs using a simple, greedy approach: ensure each triangle has edge rays forming a consistent orientation (e.g., clockwise), growing regions of alternating triangle ray orientations where possible. For remaining triangles, if there are already three edge rays, we choose the ray orientation matching two of the rays. Otherwise we set a random ray orientation. This takes 30 ms to compute for an example 40k triangle mesh, resulting in only $\sim$10of edges being traced in both directions.
@@ IV Experiments and Results
In this section, we evaluate the volume coverage accuracy of various collision body representations as well as the computational performance and collision-detection accuracy of the proposed algorithms compared to other approaches, with the Flexible Collision Library (FCL) [10] as a CPU baseline and the cuRobo library [9] providing comparisons to state-of-the-art GPU-based methods. We did not compare to Mochi [8] as the library is not open-sourced and adds unnecessary complexity for our application (see Sec. V-A1).
@@ IV-A Robot and Collision Environments
To benchmark computational speed and collision-checking accuracy of our method, we constructed several versions of a collision scene, adding obstacle meshes [26] to increase complexity as shown in Fig. 2. The simple, medium, and dense versions of the scene consisted of over 15k / 22k, 71k / 107k, and 191k / 321k triangles / edges, respectively. We performed all experiments on an AMD Ryzen 7 5700X CPU and NVIDIA RTX 3070 GPU, using g++13, CUDA 12.4, and OptiX 7.7, and we pre-allocated all GPU memory buffers avoid dynamic allocations on critical compute paths.
@@ IV-B Volume-Coverage Accuracy Measurement
Various works use different volume representations (implicit or explicit) of discrete robot poses and continuous (swept) robot pose paths to perform boolean collision detection. Though collision detection is the end objective, it is difficult to get a holistic measurement of the accuracy of these volume representations through just collision detection experiments as the results depend heavily on the set of obstacles being checked against.
@@
To generate discrete test robot poses, we used Halton sampling [30] in C-space. To generate test pose trajectories, we then sampled 30 trajectories of varying lengths connecting Halton-sampled joint configuration pairs via straight lines in C-space. The resulting volume accuracy results are detailed in Table I. Detailed cubic spline results are omitted as they showed minimal accuracy improvement over the quadratic spline while significantly increasing runtime.
@@ IV-C Discrete Pose Collision Detection Evaluation
We tested our discrete-pose collision detection methods on a set of 4096 Halton-sampled robot poses, checking against the three versions of the scene as described in Sec. IV-A. To measure the effect of batch sizes on speed, we tested our methods and cuRobo’s on power-of-2 batch sizes up to 4096 (we measured the FCL CPU baseline at a batch size of 1).
Fig. 4 shows speed results, including separate results for the ObsToRobot/RobotToObs RT-DCD methods and both combined for the highest accuracy (two-way). Despite performing a much more complex collision check than cuRobo (mesh-to-mesh vs. mesh-to-sphere), our methods outpaced cuRobo by up to 3x on medium to dense scenes, scaling past CPU performance at smaller batch sizes. The cuRobo method catches up to our two-way method at thousands of poses per batch, but such large batch sizes are often not needed in motion planning (e.g., PRM can use one large batch, but most planners use multiple waves of smaller batches).
The cuRobo method does perform faster in simple scenes. This is because (1) our mesh-based methods check many more primitives ($>$35k robot edge rays vs. 62 spheres), and (2) the hardware-accelerated BVH traversal we use scales better with denser meshes (compared to the warp library underlying cuRobo, which using a single normal CUDA thread for each point-to-mesh BVH traversal). Due to (1) and detailed mesh representations, our methods can have a higher startup cost, which can easily be reduced by simplifying the meshes rays are traced along. However, as the meshes get more detailed, the scaling of hardware-accelerated BVH traversal leads to higher speeds.

### v2

# Note Regarding ICRA Publication

@@ Abstract
This paper presents a set of simple and intuitive robot collision detection algorithms that show substantial scaling improvements for high geometric complexity and large numbers of collision queries by leveraging hardware-accelerated ray tracing on GPUs. It is the first leveraging hardware-accelerated ray-tracing for direct volume mesh-to-mesh discrete collision detection and applying it to continuous collision detection. We introduce two methods: Ray-Traced Discrete-Pose Collision Detection for exact robot mesh to obstacle mesh collision detection, and Ray-Traced Continuous Collision Detection for robot sphere representation to obstacle mesh swept collision detection, using piecewise-linear or quadratic B-splines. For robot link meshes totaling 24k triangles and obstacle meshes of over 190k triangles, our methods were up to 2.8 times faster in batched discrete-pose queries than a state-of-the-art GPU-based method using a sphere robot representation. For the same obstacle mesh scene, our sphere-robot continuous collision detection was up to 7 times faster depending on trajectory batch size. We also performed detailed measurements of the volume coverage accuracy of various sphere/mesh pose/path representations to provide insight into the tradeoffs between speed and accuracy of different robot collision detection methods.
@@ I Introduction
The second scaling issue is the large number of collision detection queries needed for planning algorithms, often on the order of thousands of pose checks in addition to sub-path/edge checks. Additionally, a robot performing complex tasks may perform hundreds of motion planning calls when evaluating candidate tasks and ways to perform each task. Parallelism from CPU multithreading or SIMD instructions (as in [5, 6]) saturates very quickly. GPUs scale much better, able to execute simultaneous instructions from thousands of times more threads. However, GPU-based collision detection [7] is difficult to implement—high-performance BVH construction and traversal on the GPU is very complex—leading to few well-maintained software implementations.
@@
    Ray-Traced Discrete-Pose Collision Detection (RT-DCD): Performs simple and intuitive mesh-to-mesh pose collision detection up to 2.8x faster than GPU-based mesh-to-sphere counterpart cuRobo [10] and up to 27x faster than CPU-based FCL [11], e.g., checking 4,096 articulated-arm poses against a >>190k triangle mesh collision scene in 4 ms.
@@
    Ray-Traced Continuous Collision Detection (RT-CCD): Constructs swept volumes of sphere-approximated robots along trajectories using OptiX curves and spheres and detects collisions with mesh obstacles. Performs up to 7x faster than cuRobo, depending on trajectory batch size and curve type (piecewise linear vs. quadratic B-spline).
@@ II Background and Related Works
Collision checking becomes more time-consuming as the geometric complexity of the robot and obstacles increases. It is quite common to simplify geometric representations of the robot and/or obstacles (e.g., using primitive shapes such as spheres or capsules), but this may be too conservative in cluttered scenes or tight spaces. Broad-phase/narrow-phase methods can also improve computational scaling through quick elimination of groups of collision geometries via different types of bounding volume hierarchies (BVH), such as AABB [1], OBB [2], SSV [3], and k-DOP [4].
@@ II-A CPU-Based DCD Methods
The Flexible Collision Library (FCL) [11] collision checks in 3D space with BVH acceleration and has seen perhaps the most widespread adoption in robotics, valued for its support for many geometry, collision pair, and query types. VAMP [5] and CAPT [6] further accelerated collision detection of sphere robots through CPU SIMD instructions. However, these libraries are CPU-bound, and CPU multithreading, SIMD, and BVH scaling are often not enough to handle both the geometric complexity and quantity of collision queries needed for low-conservatism real-time planning.
@@ II-B GPU-Based DCD Methods without Ray Tracing
Some recent methods like PairwiseNet [15] and ClearanceNet [16] leverage neural networks and GPUs for higher throughput, but ClearanceNet only achieved 91-96% detection accuracy and PairwiseNet requires training on each mesh pair for collision detection, whereas our DCD method detects all penetrating collisions between arbitrary meshes with minimal precomputation time.
@@ II-C CCD Methods
Conservative advancement (CA) methods [21, 22, 23] offer an alternative by computing a motion bound to determine the maximum forward sweep time step on a trajectory. The cuRobo library [10] uses a modified bisection and CA method on robots approximated by spheres, achieving CCD with excellent batch performance through parallel execution on GPUs. Our research proposes a new CCD method for sphere-approximated robots, using a GPU-based ray-tracing method to check against explicitly constructed swept volumes and achieving faster speeds depending on trajectory batch size and higher accuracy through swept curves which are higher order than cuRobo’s method allows.
@@ II-D Hardware-Accelerated Ray Tracing
We have previously explored hardware-accelerated ray tracing for computing signed distances (which can serve as collision detection), finding contact points, and formulating trajectory optimization constraints between robot and obstacle meshes [8]. The Mochi engine [9] explored applications to DCD between spheres, triangle meshes, and generic geometries which can be AABB-bounded and have a point-containment function. However, its mesh DCD does not treat meshes as volumes (it cannot detect an obstacle fully contained in a robot mesh or vice versa), and it has suboptimalities for common robot DCD applications (for example, not distinguishing desired robot/obstacle collision pairs and not exploring the best batching strategies for multi-query multi-link robot applications).
@@ III-A Ray Tracing Volume Mesh-to-Mesh Collision Detection
A further corner case is when the robot is fully contained inside an obstacle or vice versa, which is common in aerial, underwater, or space motion planning scenes with unanchored robots / obstacles. For these cases, it is sufficient to have watertight meshes with correctly oriented triangles (normals facing out) and to trace one maximum-length ray from the interior of each mesh toward all target meshes. If the ray hits more triangle back faces than front (which can be detected through OptiX), the ray origin is in collision. Note that this also handles target meshes that overlap.
These tracing strategies produce exact volume mesh-to-mesh detection of penetrating collisions, which is sufficient for most robot motion planning applications.
@@ III-B Discrete-Pose Collision Detection
We developed three approaches to detect collisions between robot and obstacle volume meshes: ObsToRobot (Fig.LABEL:fig:_headline_d) traces rays along obstacles toward robot links, RobotToObs (Fig.LABEL:fig:_headline_r) traces rays along robot links posed by forward kinematics toward obstacles, and “two-way” simply combines the results of ObsToRobot and RobotToObs for exact detection results. Alg. 1 shows the inputs of each, including a robot kinematics / link mesh model and a mesh obstacle scene. The output is a boolean array indicating whether each robot configuration is in collision with an obstacle. Each algorithm is GPU-parallelized to batch operations such as forward kinematics across robot configurations and aggregating results across rays, and each uses the method in Sec. III-A to identify in-collision mesh pairs.
@@
[…]
@@
To increase speedup, we precompute an oriented bounding box (OBB) for each obstacle and robot mesh via a simple PCA method taking $\sim$300us for a mesh with 20,000 vertices [25]. We then use these OBBs for broad-phase collision detection (i.e. single-level BVHs for simplicity), eliminating a large number of unnecessary posed links from IASs and rays from tracing steps. For example, in the scene shown in Fig 2, only 1/4th of the associated OBBs are in collision. To aggregate useful posed links and rays from broadphase result masks, we use stream compaction [26] a parallel algorithm that gathers subsets of data of interest. Although stream compaction has poor scaling compared other common classes of parallel algorithms on the GPU, we found it was more than worth it to reduce inputs to ray-tracing steps, which tend to take by far the majority of compute time.
@@ III-C Continuous Collision Detection
For piecewise-linear swept volumes, we use batch forward kinematics to compute robot link sphere centers on poses along a path, linearly interpolating sphere centers between subsequent poses. For more accurate path approximations, we also implemented quadratic and cubic B-splines [27], for which we compute control points from path waypoints via least squares as in [28].
Since the number of control points determines the least-squares pseudo-inverse matrix, each control point becomes a simple constant-weighted sum of all trajectory points. Thus we store this pseudo-inverse matrix in GPU constant memory for efficient broadcasting across instances of a parallel reduction kernel used to compute the curve control points across all links and pose paths.
One challenge is that OptiX meshes, spheres, and curves are all considered hollow, but for curves (and spheres starting in OptiX 9.0), ray traces against them cannot detect back-face hits (intersections starting from a curve’s interior). To address this, we choose ray directions along the edges of each obstacle mesh such that they form a strongly-connected directed graph. This guarantees catching vertex collisions as any out-of-collision vertex must have a ray-directed path to any in-collision vertex. We precompute these graphs using a simple, greedy approach: ensure each triangle has edge rays forming a consistent orientation (e.g., clockwise), first by growing regions of alternating triangle orientations where possible. For remaining triangles, if there are already three edge rays, we choose the orientation matching two of the rays. Otherwise we set a random orientation. This takes 30 ms to compute for an example 40k triangle mesh, resulting in only $\sim$10% of edges being traced in both directions.
@@ IV Experiments and Results
In this section, we evaluate the volume coverage accuracy of various collision body representations as well as the computational performance and collision-detection accuracy of the proposed algorithms compared to other approaches, with the Flexible Collision Library (FCL) [11] as a CPU baseline and the cuRobo library (with CUDA graphs enabled and self-collision disabled) [10] providing comparisons to state-of-the-art GPU-based methods. We did not compare to Mochi [9] as the library is not open-sourced and adds unnecessary complexity for our application (see Sec. V-A1).
@@ IV-A Robot and Collision Environments
To benchmark computational speed and collision-checking accuracy of our method, we constructed several versions of a collision scene, adding obstacle meshes [29] to increase complexity as shown in Fig. 2. The simple, medium, and dense versions of the scene consisted of over 15k / 22k, 71k / 107k, and 191k / 321k triangles / edges, respectively. We performed all experiments on an AMD Ryzen 7 7700x CPU and an RTX 4070 Super GPU. We wrote the library in C++/CUDA with OptiX 8.0, compiling with g++13 and CUDA 12.6, and we pre-allocated all GPU memory buffers to avoid dynamic allocations on critical compute paths.
@@ IV-B Volume-Coverage Accuracy Measurement
Various works use different volume representations (implicit or explicit) of discrete robot poses and continuous (swept) robot pose paths to perform boolean collision detection. Though collision detection is the end objective, it is difficult to get a holistic measurement of the accuracy of these volume representations through just collision detection experiments as the results depend heavily on the selected poses and set of obstacles being checked against.
@@
To generate discrete test robot poses, we used quasi-random Halton sampling [33] in C-space. To generate test pose trajectories, we then sampled 30 trajectories of varying lengths connecting Halton-sampled joint configuration pairs via straight lines in C-space. The resulting volume accuracy results are detailed in Table I. Cubic spline results are omitted as they showed minimal accuracy improvement over the quadratic spline while significantly increasing runtime.
@@ IV-C Discrete Pose Collision Detection Evaluation
We tested our discrete-pose collision detection methods on a set of 4096 Halton-sampled robot poses, checking against the three versions of the scene as described in Sec. IV-A. To measure the effect of batch sizes on speed, we tested our methods and cuRobo’s on power-of-2 batch sizes up to 4096 (we measured the FCL CPU baseline at a batch size of 1). We included all steps of Alg. 1 in benchmarking, including forward kinematics on the GPU.
Fig. 4 shows speed results, including separate results for the ObsToRobot/RobotToObs RT-DCD methods and both combined for the highest accuracy (two-way). Despite performing a much more complex collision check than cuRobo (mesh-to-mesh vs. mesh-to-sphere), our methods outpaced cuRobo by up to 2.8x on medium to dense scenes, scaling past CPU performance at smaller batch sizes. The cuRobo method catches up to our two-way method at thousands of poses per batch, but such large batch sizes are often not needed in motion planning (e.g., PRM can use one large batch, but most planners use waves of smaller batches).
The cuRobo method does perform faster in simple scenes. This is because (1) our mesh-based methods check many more primitives (up to >>35k robot edge rays even on the simple scene vs. 62 spheres), and (2) the hardware-accelerated BVH traversal we use scales better with denser meshes (compared to the warp library underlying cuRobo, which using a single normal CUDA thread for each point-to-mesh BVH traversal). The cost of (1) is not sufficiently diminished by our simple broadphase method or offset by (2) on “simple” scenes. This can be addressed through improved broadphase or conservative simplification of the meshes rays are traced along. However, as the meshes traced against get more detailed, the scaling of hardware-accelerated BVH traversal do lead to higher speeds overall.